OpenStudy (goformit100):

A natural number ends in 2. If we move this digit 2 to the beginning of the number, then the number will be doubled. Find the smallest number with this property.

OpenStudy (goformit100):

@Numb3r1

OpenStudy (anonymous):

ab...2=2ab.../2 more specifically, whatever number ab... is (multiple digits, clearly, as 2d cannot equal d2*2 because that would be (2d)4, but 42 is not twice 24 or vice versa) ab... times ten plus two equals 2*10^(?) plus ab....

OpenStudy (goformit100):

ok

OpenStudy (anonymous):

That means 2*10^(?)=9(ab...) plus 2. ab... must end in a 2, because 9*ab.... ends in an 8 to make 9*ab...+2 end in a zero.

OpenStudy (goformit100):

ok

OpenStudy (anonymous):

Wait, that has so many issues... ignore it.

OpenStudy (goformit100):

ok

OpenStudy (anonymous):

Here's a fixed version:

OpenStudy (goformit100):

ok

OpenStudy (anonymous):

That means 2*10^(?)+ab...=20(ab...) plus 2. The twenty accounts for the doubling of the original number.

OpenStudy (anonymous):

Therefore, 2*10^(?)=19(ab...) plus 2. 2*10^(?)-2 looks like 19....98. Divide by 19 and we have 1...?

OpenStudy (anonymous):

Nope, that makes no sense.

OpenStudy (anonymous):

ab2=2ab/2... let's see... ab must start in a 1, so I was right anyway.

OpenStudy (anonymous):

Unless we carry... that's it!

OpenStudy (goformit100):

ok

OpenStudy (anonymous):

2ab2=ab22*2... nope. I should just go to sleep. Good luck!

OpenStudy (goformit100):

ok

OpenStudy (goformit100):

Thank you sir

OpenStudy (anonymous):

105263157894736842